OpenStudy (anonymous):

need to solve by graphing 12=2xsquared -5x ? help

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

it gives me answers but idk how to work it out a. x=8,-3 b.x=-8,3 c.x=-3/2,4 d.x=3,-4

OpenStudy (anonymous):

2x^2-5x-12=0 so find the roots of this and u'll get answer as A